OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that passes through the given point and its parallel to the graph for each equation.

OpenStudy (anonymous):

(1,-4) 9x+3y=8 M=9 Point: (-1,-4) Point slope: y-y1= m(x-x1) y+4=9x+9 - 4 -4 y=5x+9 b=9 m= 5/1 Is this right?